# Bifurcation sequences in the secular 3D planetary 3-Body problem: a geometric approach

### v2

@@ 3.2 Hopf representation of the phase space
$\sigma_{1}=X_{2}X_{3}+Y_{2}Y_{3}\,,\quad\sigma_{2}=Y_{2}X_{3}-Y_{3}X_{2}\,,\quad\sigma_{3}=\frac{1}{2}\left(X_{2}^{2}+Y_{2}^{2}-X_{3}^{2}-Y_{3}^{2}\right)\leavevmode\nobreak\ ,$ (24)
@@
$\sigma_{3}=W_{2}-W_{3}=2\Gamma\,.$
We also have the relation $\sigma_{0}=W_{2}+W_{3}=2J$, as well as
@@
Given the values of $(\sigma_{1},\sigma_{2},\sigma_{3})$, the values of $\Gamma$, $J$ and $\psi$ can be computed unequivocally using the relations (26) and (27). Furthermore, the Hamiltonian $\mathcal{Z}$ is even in the difference $w_{2}-w_{3}$, hence it does not depend on $\sigma_{2}$, i.e., $\mathcal{Z}=\mathcal{Z}(\sigma_{1},\sigma_{3};\sigma_{0})$. Fixing a value of $\sigma_{0}$ (i.e. of the integral $J=\sigma_{0}/2$), the continuous in time phase flow under $\mathcal{Z}$ can be expressed in the Hopf variables via the equations
$\dot{\sigma}_{1}=\{\sigma_{1},\sigma_{3}\}{\partial\mathcal{Z}\over\partial\sigma_{3}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak$ $\dot{\sigma}_{2}=\{\sigma_{2},\sigma_{1}\}{\partial\mathcal{Z}\over\partial\sigma_{1}}+\{\sigma_{2},\sigma_{3}\}{\partial\mathcal{Z}\over\partial\sigma_{3}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \dot{\sigma}_{3}=-\{\sigma_{1},\sigma_{3}\}{\partial\mathcal{Z}\over\partial\sigma_{1}}\leavevmode\nobreak\ \leavevmode\nobreak\ .$ (28)
In particular, the level curves of Fig. 2 can be represented by the intersection of the constant energy surface $\mathcal{Z}(\sigma_{1},\sigma_{3};\sigma_{0})=\mathcal{E}$ with the sphere (27). Altering the energy $\mathcal{E}$ in the interval $\mathcal{E}_{L}(\sigma_{0})\leq\mathcal{E}\leq\mathcal{E}_{R}(\sigma_{0})$, while keeping $\sigma_{0}=2J$ constant yields a family of closed curves on the sphere which can be mapped to the invariant curves of the Poincaré surface of section of Fig. 2 through the relations $X_{2}=-\sqrt{2(\Gamma+J)}\cos(\psi-\pi)$, $Y_{2}=\sqrt{2(\Gamma+J)}\sin(\psi-\pi)$, with $\Gamma=\sigma_{3}/2$, $\cos\psi=\sigma_{1}/(\sqrt{\sigma_{0}^{2}-\sigma_{3}^{2}})$, $\sin\psi=\sigma_{2}/(\sqrt{\sigma_{0}^{2}-\sigma_{3}^{2}})$. Equivalently, keeping the energy $\mathcal{E}$ fixed and altering $\sigma_{0}$ in the interval $\sigma_{0,min}(\mathcal{E})\leq\sigma_{0}\leq\sigma_{0,max}(\mathcal{E})$ allows to recover the family of all invariant curves in Fig. 2.
@@ 3.3 Geometrical representation of the sequence of bifurcations
Consider a fixed integral approximation $\mathcal{Z}$ to the Hamiltonian $\mathcal{H}_{sec}$. Let
@@
The whole sequence of bifurcations of new fixed points as in Fig. 2 can be computed exploiting the above properties as follows: since the Hamiltonian does not depend on $\sigma_{2}$, for any fixed value of $\sigma_{0}$ all possible kinds of intersection or tangency of the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ with $\mathcal{S}_{\sigma_{0}}$ can be classified through their projection to the plane $\sigma_{2}=0$. In particular, all points of tangency are along the meridian of the sphere $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ contained in the plane $(\sigma_{1},\sigma_{3})$ with $\sigma_{2}=0$. Altering the energy in the interval $\mathcal{E}_{L}(\sigma_{0})\leq\mathcal{E}\leq\mathcal{E}_{R}(\sigma_{0})$ we then compute the curve $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$ at which the surface $\mathcal{C}_{\sigma_{0},\mathcal{E}}$ intersects the same plane, for values of $\mathcal{E}$ within the permissible limits discussed in the previous subsection. We then distinguish two cases, both satisfying the equation
$\nabla{\mathcal{\mathcal{Z}}=\mu\,(\sigma_{1},\sigma_{2},\sigma_{3})}$ (31)
@@
$\sigma_{2}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \nabla_{\sigma_{1},\sigma_{3}}\left(\mathcal{Z}\right)_{\sigma_{2}=0}\neq 0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{rank}\begin{pmatrix}\displaystyle{2\sigma_{1}}&\displaystyle{2\sigma_{3}}\\ \displaystyle{\frac{\partial\mathcal{Z}}{\partial\sigma_{1}}}&\displaystyle{\frac{\partial\mathcal{Z}}{\partial\sigma_{3}}}\end{pmatrix}=1\leavevmode\nobreak\ \leavevmode\nobreak\ ,$
@@
$\sigma_{2}=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \sigma_{3}\frac{\partial\mathcal{Z}}{\partial\sigma_{1}}=\sigma_{1}\frac{\partial\mathcal{Z}}{\partial\sigma_{3}}\leavevmode\nobreak\ .$ (32)
@@
The critical points (CP) of tangency are called of the ‘first kind’ (CP1), according to the terminology introduced in [7], and they represent periodic orbits under the complete 2-DOF flow of the Hamiltonian (23) (see subsection 3.1). We distinguish two subcases (see Fig. 4):
Figure 4: An example of inner tangency between $\mathcal{S}_{\sigma_{0}}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{1})}}$ at the point $P_{1}$ (left), and outer tangency at the point $P_{2}$ (right). The example is taken from the analysis of the integrable model $\mathcal{Z}=\mathcal{H}_{int}$ (see second panel of the second row of Fig. 6, corresponding to the phase portrait $\mathcal{P_{Z}}(\sigma_{0})$ with $=6.15\cdot 10^{-3})$).
Inner tangency: in a neighborhood enclosing the point of tangency, the curve $\mathcal{C}_{\sigma_{0},\mathcal{E}^{(P_{1})}}^{(\sigma_{2}=0)}$ is contained within the disc limited by the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ (point $P_{1}$ in Fig. 4). Then, in the same neighborhood the constant energy surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{1})}}$ intersects the sphere $\mathcal{S}_{\sigma_{0}}$ at two closed loops joining each other at the singular point $P_{1}$. The fixed point $P_{1}$ corresponds to a Floquet-unstable periodic orbit, and the two loops to the separatrices asymptotically connected to the periodic orbit.
Outer tangency: in a neighborhood enclosing the point of tangency, the curve $\mathcal{C}_{\sigma_{0},\mathcal{E}^{(P_{2})}}^{(\sigma_{2}=0)}$ is not contained within the disc limited by the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ (point $P_{2}$ in Fig. 4). Then, the constant energy surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{2})}}$ has no common points with the sphere $\mathcal{S}_{\sigma_{0}}$ in the same neighborhood other than $P_{2}$. The fixed point corresponds to a Floquet-stable periodic orbit. For values of the energy $\mathcal{E}\approx\mathcal{E}^{(P_{2})}$ we obtain surfaces $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ which intersect the sphere at closed curves near to, and surrounding $P_{2}$. These are invariant curves of the reduced flow representing 2-tori of the complete flow around the stable periodic orbit.
@@
Figure 5: Graphical representation of the critical points of second kind $F_{1}$ and $F_{2}$. Top: elliptic case (see text). The example is taken from the choice of integrable model $\mathcal{Z}=\mathcal{H}_{int}$, and it corresponds to the phase portrait $\mathcal{P_{Z}}(\sigma_{0})$ with $\sigma_{0}=5.93\cdot 10^{-3}$ (the third panel in the top row of Fig. 6). Bottom: hyperbolic case. The example is taken from the choice of integrable model $\mathcal{Z}=\mathcal{H}_{int}^{(1)}$, and it corresponds to the phase portrait $\mathcal{P_{Z}}(\sigma_{0})$ with $\sigma_{0}=8.94\cdot 10^{-3}$ (see Fig. 8). In both cases, the left panel shows the intersection of the sphere $\mathcal{S}_{\sigma_{0}}$ with an energy surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ for $\mathcal{E}\approx\mathcal{E}^{(F)}$ and ~$\mathcal{E}\approx\tilde{\mathcal{E}}^{(F)}$, the middle panel for $\mathcal{E}=\mathcal{E}^{(F)}$ and =~$\mathcal{E}=\tilde{\mathcal{E}}_{1}^{(F)}$, and the right panels the corresponding projections to the plane $\sigma_{2}=0$ (intersection of the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ with the circles $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$). We have $\mathcal{E}=\mathcal{E}_{1}^{(F)}=-1.7362\cdot 10^{-5}$ and $\mathcal{E}^{(F)}=-1.7366\cdot 10^{-5}$ in the top row, in which the corresponding surfaces are a cylinder (left panel) reducing to a straight line (middle panel). We have =~$\mathcal{E}=\tilde{\mathcal{E}}_{1}^{(F)}=-2.874\cdot 10^{-5}$ and ~$\tilde{\mathcal{E}}^{(F)}=-2.87453\cdot 10^{-5}$ in the bottom panel, where the corresponding surfaces are hyperbolic sheets (left panel) reducing to two sheets intersecting at the line $F_{1}F_{2}$ in the middle panel.
@@
(that is, Eq. (31) with $\mu=0$) the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(F)}}$, with $\mathcal{E}^{(F)}=\mathcal{Z}(\sigma_{1,F},\sigma_{3,F};\sigma_{0})$ reduces either to a single straight line parallel to the $\sigma_{2}-$axis, or to two sheets intersecting along a line parallel to the $\sigma_{2}-$axis. If the line intersects transversally the sphere $\mathcal{S}_{\sigma_{0}}$ (see Fig. 5) we obtain two fixed points of the flow $F_{1},F_{2}$ on the sphere, i.e. two periodic orbits of the complete flow. Such critical points (or orbits) are called of the ‘second kind’ (CP2). Their stability depends on the sign-definiteness of the quadratic form $(\partial^{2}\mathcal{Z}/\partial\sigma_{1}^{2})\delta\sigma_{1}^{2}+2(\partial^{2}\mathcal{Z}/\partial\sigma_{1}\partial\sigma_{3})\delta\sigma_{1}\delta\sigma_{3}+(\partial^{2}\mathcal{Z}/\partial\sigma_{3}^{2})\delta\sigma_{3}^{2}$ where the partial derivatives are evaluated at $\sigma_{1}=\sigma_{1,F}$, $\sigma_{3}=\sigma_{3,F}$. If the quadratic form is sign-definite, the points $F_{1},F_{2}$ are linearly stable. For values of the energy $\mathcal{E}$ near $\mathcal{E}^{(F)}$ the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ has the form of a cylinder with elliptic cross-section which surrounds the axis $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(F)}}$, and its intersections with the sphere are elliptic-like curves which surround the fixed points $F_{1}$ or $F_{2}$ and represent 2-tori of the complete flow around the periodic orbits corresponding to the points $F_{1}$ or $F_{2}$ (see top row in Fig. 5). In the case, however, where the sign of the quadratic form is indefinite, the points $F_{1},F_{2}$ represent unstable periodic orbits, and the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(F)}}$, together with nearby surfaces for energies close to $\mathcal{E}^{(F)}$ intersect the sphere forming saddles around each of the points $F_{1}$ or $F_{2}$ (bottom row in Fig. 5).

### v3

@@ 3.2 Hopf representation of the phase space
$\sigma_{1}=X_{2}X_{3}+Y_{2}Y_{3}\,,\quad\sigma_{2}=Y_{2}X_{3}-Y_{3}X_{2}\,,\quad\sigma_{3}=\frac{1}{2}\left(X_{2}^{2}+Y_{2}^{2}-X_{3}^{2}-Y_{3}^{2}\right)\penalty 10000\ ,$ (28)
@@
$\sigma_{3}=W_{2}-W_{3}=2\Gamma\,$,
where the variables $(\psi,\varphi,\Gamma,J)$ are introduced in Eq. (26). We also have the relation $\sigma_{0}=W_{2}+W_{3}=2J$, as well as
@@
Given the values of $(\sigma_{1},\sigma_{2},\sigma_{3})$, the values of $\Gamma$, $J$ and $\psi$ can be computed unequivocally using the relations (30) and (31). Furthermore, the Hamiltonian $\mathcal{Z}$ is even in the difference $w_{2}-w_{3}$, hence it does not depend on $\sigma_{2}$, i.e., $\mathcal{Z}=\mathcal{Z}(\sigma_{1},\sigma_{3};\sigma_{0})$. Fixing a value of $\sigma_{0}$ (i.e. of the integral $J=\sigma_{0}/2$), the phase flow under $\mathcal{Z}$ , continuous in time, can be expressed in the Hopf variables via the equations
$\dot{\sigma}_{1}=\{\sigma_{1},\sigma_{3}\}{\partial\mathcal{Z}\over\partial\sigma_{3}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\$$\dot{\sigma}_{2}=\{\sigma_{2},\sigma_{1}\}{\partial\mathcal{Z}\over\partial\sigma_{1}}+\{\sigma_{2},\sigma_{3}\}{\partial\mathcal{Z}\over\partial\sigma_{3}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \dot{\sigma}_{3}=-\{\sigma_{1},\sigma_{3}\}{\partial\mathcal{Z}\over\partial\sigma_{1}}\penalty 10000\ \penalty 10000\ .$ (32)
In particular, the level curves of Fig. 3 can be represented by the intersection of the constant energy surface $\mathcal{Z}(\sigma_{1},\sigma_{3};\sigma_{0})=\mathcal{E}$ with the sphere (31). Altering the energy $\mathcal{E}$ in the interval $\mathcal{E}_{L}(\sigma_{0})\leq\mathcal{E}\leq\mathcal{E}_{R}(\sigma_{0})$, while keeping $\sigma_{0}=2J$ constant yields a family of closed curves on the sphere which can be mapped to the invariant curves of the Poincaré surface of section of Fig. 3 through the relations $X_{2}=-\sqrt{2(\Gamma+J)}\cos(\psi-\pi)$, $Y_{2}=\sqrt{2(\Gamma+J)}\sin(\psi-\pi)$, with $\Gamma=\sigma_{3}/2$, $\cos\psi=\sigma_{1}/\sqrt{\sigma_{0}^{2}-\sigma_{3}^{2}}$, $\sin\psi=\sigma_{2}/\sqrt{\sigma_{0}^{2}-\sigma_{3}^{2}}$. Equivalently, keeping the energy $\mathcal{E}$ fixed and altering $\sigma_{0}$ in the interval $\sigma_{0,min}(\mathcal{E})\leq\sigma_{0}\leq\sigma_{0,max}(\mathcal{E})$ allows to recover the family of all invariant curves in Fig. 3.
The special motions leading to singularities in ref. [21] can be seen in the Hopf variables in the following way. Concerning i) circular trajectories: the inner circular trajectory $\mathrm{e}_{2}=0$ is described by a point in the sphere, more precisely by the south pole $(\sigma^{(S)}_{0},\sigma^{(S)}_{1}=0,\sigma^{(S)}_{2}=0,\sigma^{(S)}_{3}=-\sigma^{(S)}_{0})$. In the Poincaré surface of section $Y_{3}=0$ $\dot{Y}_{3}\geq 0$ this would lead to a curve passing through the origin. Instead, the outher circular trajectory $\mathrm{e}_{3}=0$ is described by a north pole in the Hopf variables, i.e. $(\sigma^{(N)}_{0},\sigma^{(N)}_{1}=0,\sigma_{2}^{(N)}=0,\sigma^{(N)}_{3}=\sigma^{(N)}_{0})$ and by an open curve in the Poincaré surface of section. This has been already analyzed in Fig. 11 of [17]. Moreover, recalling the expression of the cosine of the mutual inclination as function of the eccentricity (see Eq. (29) of [17]) and considering that it is in [1,1][-1,1], it is possible to find the limits for such a kind of orbits. On the other hand, the ii) coplanar motions can be limited by the fixed value of the $\mathrm{AMD}$ (see Fig. 8 of [17], where the existence of coplanar orbits in a given interval of energies is provided).
@@ 3.3 Geometrical representation of the sequence of bifurcations
Consider a fixed integrable approximation $\mathcal{Z}$ to the Hamiltonian $\mathcal{H}_{sec}$. Let
@@
The whole sequence of bifurcations of new fixed points as in Fig. 3 can be computed exploiting the above properties as follows: since the Hamiltonian does not depend on $\sigma_{2}$, for any fixed value of $\sigma_{0}$ all possible kinds of intersection or tangency of the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ with $\mathcal{S}_{\sigma_{0}}$ can be classified through their projection to the plane $\sigma_{2}=0$. In particular, all points of tangency are along the meridian of the sphere $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ contained in the plane $(\sigma_{1},\sigma_{3})$ with $\sigma_{2}=0$. Altering the energy in the interval $\mathcal{E}_{L}(\sigma_{0})\leq\mathcal{E}\leq\mathcal{E}_{R}(\sigma_{0})$ we then compute the curve $\mathcal{C}_{\sigma_{0},\mathcal{E}}^{(\sigma_{2}=0)}$ at which the surface $\mathcal{C}_{\sigma_{0},\mathcal{E}}$ intersects the plane $\sigma_{2}=0$, for values of $\mathcal{E}$ within the permissible limits discussed in the previous subsection. We then distinguish two cases, both satisfying the equation
$\nabla\mathcal{\mathcal{Z}}={2}\mu\,\begin{pmatrix}\sigma_{1}\\ \sigma_{2}\\ \sigma_{3}\end{pmatrix}$ (35)
@@
$\sigma_{2}=0,\penalty 10000\ \penalty 10000\ \nabla_{\sigma_{1},\sigma_{3}}\left(\mathcal{Z}\right)_{\sigma_{2}=0}\neq 0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathrm{rank}\begin{pmatrix}\displaystyle{2\sigma_{1}}&\displaystyle{2\sigma_{3}}\\ \displaystyle{\frac{\partial\mathcal{Z}}{\partial\sigma_{1}}}&\displaystyle{\frac{\partial\mathcal{Z}}{\partial\sigma_{3}}}\end{pmatrix}=1\penalty 10000\ \penalty 10000\ ,$
@@
$\sigma_{2}=0,\penalty 10000\ \penalty 10000\ \penalty 10000\ \sigma_{3}\frac{\partial\mathcal{Z}}{\partial\sigma_{1}}=\sigma_{1}\frac{\partial\mathcal{Z}}{\partial\sigma_{3}}\penalty 10000\ .$ (36)
@@
The critical points (CP) of tangency are called of the ‘first kind’ ( CPI), according to the terminology introduced in [9], and they represent periodic orbits under the complete 2-DOF flow of the Hamiltonian (27) (see subsection 3.1). We distinguish two subcases (see Fig. 4):
Figure 4: An example of inner tangency between $\mathcal{S}_{\sigma_{0}}$ and $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{1})}}$ at the point $P_{1}$ (left), and outer tangency at the point $P_{2}$ (right). The example is taken from the analysis of the integrable model $\mathcal{Z}=\mathcal{H}_{int}$ (see second panel of the second row of Fig. 7, corresponding to the phase portrait $\mathcal{P_{Z}}(\sigma_{0})$ with ${\sigma_{0}}=6.15\cdot 10^{-3}$).
Inner tangency: in the space of variables $(\sigma_{1},\sigma_{2},\sigma_{3})$, there exists a neighborhood enclosing the point of tangency, where the curve $\mathcal{C}_{\sigma_{0},\mathcal{E}^{(P_{1})}}^{(\sigma_{2}=0)}$ is contained within the disc limited by the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ (point $P_{1}$ in Fig. 4). Then, in the same neighborhood the constant energy surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{1})}}$ intersects the sphere $\mathcal{S}_{\sigma_{0}}$ at two closed loops joining each other at the singular point $P_{1}$. The fixed point $P_{1}$ corresponds to a Floquet-unstable periodic orbit, and the two loops to the separatrices asymptotically connected to the periodic orbit.
Outer tangency: there exists a neighborhood enclosing the point of tangency where the curve $\mathcal{C}_{\sigma_{0},\mathcal{E}^{(P_{2})}}^{(\sigma_{2}=0)}$ is not contained within the disc limited by the circle $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$ (point $P_{2}$ in Fig. 4). Then, the constant energy surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(P_{2})}}$ has no common points with the sphere $\mathcal{S}_{\sigma_{0}}$ in the same neighborhood other than $P_{2}$. The fixed point corresponds to a Floquet-stable periodic orbit. For values of the energy $\mathcal{E}\approx\mathcal{E}^{(P_{2})}$ we obtain surfaces $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ which intersect the sphere at closed curves near to, and surrounding $P_{2}$. These are invariant curves of the reduced flow representing 2-tori of the complete flow around the stable periodic orbit.
@@
Figure 5: Graphical representation of the critical points of second kind $F_{1}$ and $F_{2}$ in the elliptic case (see text). The example is taken from the choice of integrable model $\mathcal{Z}=\mathcal{H}_{int}$, and it corresponds to the phase portrait $\mathcal{P_{Z}}(\sigma_{0})$ with $\sigma_{0}=5.93\cdot 10^{-3}$ (the third panel in the top row of Fig. 7). The top left panel shows the intersection of the sphere $\mathcal{S}_{\sigma_{0}}$ with an energy surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ for $\mathcal{E}\approx\mathcal{E}^{(F)}$ and the right top panel for $\mathcal{E}=\mathcal{E}^{(F)}$. The bottom panels show the corresponding projections to the plane $\sigma_{2}=0$ (intersection of the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ with the circles $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$). We have $\mathcal{E}=\mathcal{E}_{1}^{(F)}=-1.7362\cdot 10^{-5}$ and $\mathcal{E}^{(F)}=-1.7366\cdot 10^{-5}$, in which the corresponding surfaces are a cylinder ( top left panel) reducing to a straight line ( top right panel). The curve $\mathcal{C}_{\mathcal{Z}}(\sigma_{0},\mathcal{E}_{1}^{(F)})$ in the corresponding phase portraits $\mathcal{P_{Z}}(\sigma_{0})$ (see third panel of first row of Fig. 7) is plotted in green.
Figure 6: Graphical representation of the critical points of second kind $F_{1}$ and $F_{2}$ in the hyperbolic case (see text). The example is taken from the choice of integrable model $\mathcal{Z}=\mathcal{H}_{int}^{(1)}$, and it corresponds to the phase portrait $\mathcal{P_{Z}}(\sigma_{0})$ with $\sigma_{0}=8.94\cdot 10^{-3}$ (see Fig. 9). The top left panel shows the intersection of the sphere $\mathcal{S}_{\sigma_{0}}$ with an energy surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ for $\mathcal{E}\approx\tilde{\mathcal{E}}^{(F)}$and the right top panel for $\mathcal{E}={\tilde{\mathcal{E}}}^{(F)}$. The bottom panels show the corresponding projections to the plane $\sigma_{2}=0$ (intersection of the curves $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}^{(\sigma_{2}=0)}$ with the circles $\mathcal{S}_{\sigma_{0}}^{(\sigma_{2}=0)}$). We have $\mathcal{E}=\tilde{\mathcal{E}}_{1}^{(F)}=-2.874\cdot 10^{-5}$ and $\tilde{\mathcal{E}}^{(F)}=-2.87453\cdot 10^{-5}$, where the corresponding surfaces are hyperbolic sheets ( top left panel) reducing to two sheets intersecting at the line $F_{1}F_{2}$ in the top right panel. The curve $\mathcal{C}_{\mathcal{Z}}(\sigma_{0},\tilde{\mathcal{E}}_{1}^{(F)})$ in the corresponding phase portraits $\mathcal{P_{Z}}(\sigma_{0})$ (see second panel of second row of Fig. 9) is plotted in green.
@@
(that is, Eq. (35) with $\mu=0$) the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(F)}}$, with $\mathcal{E}^{(F)}=\mathcal{Z}(\sigma_{1,F},\sigma_{3,F};\sigma_{0})$ reduces either to a single straight line parallel to the $\sigma_{2}$-axis, or to two sheets intersecting along a line parallel to the $\sigma_{2}$-axis. If the line intersects transversely the sphere $\mathcal{S}_{\sigma_{0}}$ (see Fig. 5 and 6) we obtain two fixed points of the flow $F_{1},F_{2}$ on the sphere, i.e. two periodic orbits of the complete flow. Such critical points (or orbits) are called of the ‘second kind’ ( CPII). Their stability depends on the sign-definiteness of the quadratic form $(\partial^{2}\mathcal{Z}/\partial\sigma_{1}^{2})\delta\sigma_{1}^{2}+2(\partial^{2}\mathcal{Z}/\partial\sigma_{1}\partial\sigma_{3})\delta\sigma_{1}\delta\sigma_{3}+(\partial^{2}\mathcal{Z}/\partial\sigma_{3}^{2})\delta\sigma_{3}^{2}$ where the partial derivatives are evaluated at $\sigma_{1}=\sigma_{1,F}$, $\sigma_{3}=\sigma_{3,F}$. If the quadratic form is sign-definite, the points $F_{1},F_{2}$ are linearly stable. For values of the energy $\mathcal{E}$ near $\mathcal{E}^{(F)}$ the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}}$ has the form of a cylinder with elliptic cross-section which surrounds the axis $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(F)}}$, and its intersections with the sphere are elliptic-like curves which surround the fixed points $F_{1}$ or $F_{2}$ and represent 2-tori of the complete flow around the periodic orbits corresponding to the points $F_{1}$ or $F_{2}$ (see Fig. 5). In the case, however, where the sign of the quadratic form is indefinite, the points $F_{1},F_{2}$ represent unstable periodic orbits, and the surface $\mathcal{C}_{\sigma_{0},\,\mathcal{E}^{(F)}}$, together with nearby surfaces for energies close to $\mathcal{E}^{(F)}$ intersect the sphere forming saddles around each of the points $F_{1}$ or $F_{2}$ (see Fig. 6).